OpenStudy (anonymous):

MEDAL AND FAN!!! Suppose the area of a rectangle was 31.5 feet. If the length and width of the rectangle were each multiplied by 0.2, what is the area of the resulting rectangle. Explain.

OpenStudy (javk):

let lengh=l and width=w Step 1: Write down an eqn to what happened \[l \times w= 31.5\] now lets say we multiply both the length and the width by 0.2 what would happen... \[(0.2\times l) \times (0.2\times w)\] lets factorise this now \[0.2^{2}(l \times w)\] going back to the above eqn. lets compare we multiplied the right hand side by 0.2^2 therefore we must multiply the left hand side by 0.2^2 too \[0.2^{2}\times (l \times w)= 0.2^{2}\times(31.5)=1.26\] so your new area will be 1.26 squared feet Oh and remember don't panic the area was meant to decrease, because by multiplying the sides by 0.2 you are effectively shrinking your rectangle, so in effect the area would also reduce